Shift Summary for Owl shift, 10/Jul/99

Miguel Barrio (mbarrio@hep.uchicago.edu)
Sat, 10 Jul 1999 07:01:16 -0500

Shift Summary for Owl shift, 10/Jul/99

On Shift: Sasha Glazov; Miguel Barrio

Hours of beam available at all intensities: 6
Hours of beam available at nominal intensity: 6
Hours of E832 data (special runs not included): 6
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:
Tried to install new version of run-control at around 1:50. New version
failed and we are restoring the old version, Sidney and Taku were
notified. Time lost ~ 25 min.
Total time lost (including kfrends and such) was about 35 minutes.
Aditional downtime due to accelerator problems.

========================================================================
Runs ended this shift:

Run 13845 - 22M events - BEAM_832

Tapes brought to FCC this shift:

QKC104 QKC105 QKC106
QKS104 QKS105 QKS106
QKK104 QKK105 QKK106
QKB036

========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:
E832 datataking
Permanent (though minor) HCC problem. HCC alarm button goes red
sporadically, almost every spill. Elliot politely screams at us once
every ten spills (on average).
Constant trigger error states (about 5 times during the shift),
kfrending fixes the problem.
Top right-hand-side beam position monitor goes bananas.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13845 334 1:50 Tapes fill up, run ends.
1:55 SG installs new version of run control,
which is supposed to initialize tapes
simultaneously.
13846 2:04 SG initializes new run and run control
initializes tapes simultaneously, but run
aborts right after tape init. SG reinstalls
old version of run control. (25 minutes lost)
13847 2:14 Start of run, back to taking data.
2:55 We check the kumquat plots and we see a weak
wire (# 43 apparently). We email Mista Shift
Tsar (obscurely known as Jason Ladue) explaining
the matters.
5:20 We get an error from the HCC system: the
top-right quadrant goes blank. Three Etot resets
(reloading) are required to fix the problem
(in the meantime we kicked JasonH outta bed).
Spills 137 thru 141 were affected by this problem.
5:51 Beam is gone... Screen anounces energy is
conserved, we are puzzled...
6:16 Beam is back in record time. Yay!

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift